# InsViE-1M: Effective Instruction-based Video Editing with Elaborate Dataset Construction

### v1

@@ 1 Introduction
However, the training data construction for instruction-based video editing (InsViE) is much more challenging than that for image editing. InstructVid2Vid [27] and EffiVED [40] create edited videos utilizing one-shot methods [34, 23], which is highly time consuming. As a result, only 24K videos with 8 frames per video of 256$\times$256 resolution are provided in  [40], whose video quantity and quality are not sufficient to train a robust video editor. InsV2V [6] synthesizes 400K editing pairs; however, the source videos are synthetic, and hence the trained models are less effective when editing real-world videos. In addition, the limited video resolution (256$\times$256) of this dataset limits the application of the trained models. As summarized in Table 1, existing InsViE datasets suffer from low-resolution, short video duration, and insufficient quantity and quality, making InsViE remain a challenging task.
@@ 2 Related Work
Training-free video editing. Many existing video editing methods rely on pre-trained large image/video models [26, 10, 20, 17, 8, 7, 18, 9] to perform DDIM [30] inversion and denoising without training on any data. FateZero [26] introduces attention blending to disentangle the editing of objects and background. To ensure motion consistency, FLATTEN [8] and RAVE [17] use optical flow and random noise shuffling to improve the editing effect. Tokenflow [10] and VidToMe [20] merge the intermediate tokens to accelerate the editing process. AnyV2V [18] and Videoshop [9] adopt pre-trained image-to-video (I2V) models [4, 39] to edit the whole video with a given edited first-frame, providing more interaction with users. However, training-free methods often yield unsatisfactory results due to the inherent gap between generation and editing tasks.
@@
Training-based video editing. Recent works [5, 6, 40, 15, 21] have started to build large video editing datasets to train video editing models. ViViD-10M [15] and GenProp [21] utilize mask and the first frame to edit the video by pre-trained large models. As initiated by the work of InstructVid2Vid [27], a class of methods [5, 6, 40] adopt the instruction-based editing approach. For more robust performance, InsV2V [6] creates a large dataset with more than 400k synthetic pairs and trains the editing model by introducing temporal layers to the model of InstructPix2Pix [5]. Since image editing models often fail to preserve temporal consistency, EffiVED [40] generates 155k training pairs, including 24k real-world source videos, to train a video editing model upon video generation model. However, fine-tuning a pre-trained video generation model requires extensive high-quality triplets (source video, edited video, instruction) to achieve sufficient instruction-based editing capability. In this paper, we construct a large video editing dataset for training instruction-based editing models by collecting high-quality source videos and designing effective filtering strategies to improve training data quality.
@@ 3.1 Triplet generation from real-world videos
We curate high-quality videos from Pexel [2] and Openvid [22], and design an elaborate two-stage pipeline to generate and screen the edited samples. As shown in Fig. 2(a), we perform first frame editing and screening in the $1^{st}$ stage, and video editing and filtering in the $2^{nd}$ stage.
@@ 3.2 Triplet generation from image editing pairs
Existing image editing datasets [5, 38] possess a variety of image pairs with diverse editing types, which can be employed to generate video editing data. We first select 150k image editing triplets from InstructPix2Pix [5] based on metrics such as CLIP similarity, CLIP directional score, and the GPT-4o score we used in Sec. 3.1. We then employ all the 10k image editing pairs in MagicBrush [38] since this dataset is manually annotated with precise editing effects. As shown in Fig. 2(b), with these image editing pairs, we first generate the source videos from the source images through I2V generation model SVD [4]. Then, we perform video editing and filtering using the same process as the $2^{nd}$ stage used in Sec. 3.1. Finally, we generate 110k filtered video editing triplets from image editing pairs.
@@ 3.3 Generate static video triplets from images
Since producing video editing data is costly, previous works [15, 40, 32] have proposed to utilize images as part of data for mixed training. However, directly training on image data may not align well with real-world videos, hindering the training of video editing models. Considering the fact that a portion of real-world videos feature only on camera movement, we translate the images into static videos with manual camera operations, including (1) zoom-in, (2) zoom-out, (3) move left, (4) move right, (5) move down, (6) move up. The edited image is generated by using CosXL [3] as in the $1^{st}$ stage of Sec. 3.1. We crop each source and edited image pair at equal intervals to create image sequence pair, which is then processed with bilinear interpolation to enhance the video smoothness. Then we apply the filtering process as in the $2^{nd}$ stage of Sec. 3.1. In total, we produce 450k triplets of static videos for training.
@@ 4 InsViE Model Training
$\mathcal{L}$ $=\mathbb{E}_{\epsilon\sim\mathcal{N}(0,1),t}\left[\|\epsilon-\epsilon_{\theta}(t,z_{t},z_{source},z_{text})\|_{2}^{2}\right],\vspace{-0.1cm}$ (1)
@@ 4.2 Training and sampling
$\quad+s_{T}\cdot\left(\epsilon_{\theta}(z_{t},z_{source},z_{text})-\epsilon_{\theta}(z_{t},\varnothing,\varnothing)\right)$
@@ 5.1 Experimental Settings
Implement details. Our InsViE model is finetuned on the CogVideoX-2B model [37] using LoRA [14] with rank of 128. The training resolution is set to $720\times 480$ by random cropping on the $1024\times 576$ resolution source/edited videos in our InsViE-1M dataset. The training is conducted on 64 NVIDIA A100 GPUs for 40k iterations. The batch-size is 128. To be specific, we train our model on Set-S1, Set-S2 and Set-S3 for 20k, 10k, 10k iterations, respectively.
Compared methods. We compare our InsViE model with several representative methods, including SD-based training-free methods FateZero [26], TokenFlow [10] and RAVE [17], SVD-based training-free method Videoshop [9], and instruction-based method InsV2V [6].
@@
Evaluation metrics. Following previous methods and the benchmark in awesome-diffusion-v2v [31], we evaluate the video editing result from three aspects. (1) Temporal consistency. We use CLIP [28] temporal score to compute the CLIP embedding differences between frames, and use optical flow EPE  [35] to assess the motion smoothness. (2) Textual alignment. We employ the CLIP text-image embedding similarity and the pick score, which represents well human perception, for assessing textual alignment. (3) Video quality. We employ DOVER [33], which is a state-of-the-art video quality assessment method trained on human-ranked video datasets, for evaluating edited video quality.
In addition, for each of the three aspects mentioned above, we provide (4) the average score of GPT-4o [1], which is introduced in the scoring procedure in our data filtering stage (ranging from 1 to 5), to more comprehensively evaluate the quality of edited videos by competing methods.
Testing dataset. As in previous methods [10, 9, 17], we collect 100 video samples from DAVIS [25], YoutubeVOS [36] and Pexels [2] as the test set. We caption the videos to generate the corresponding instructions and paired captions using large VLMs [13] for instruction-based and training-free methods, respectively. Considering the fact that our method can process $720\times 480$ resolution video, while the competing methods can only process $512\times 512$ or lower resolution video, we collect 50 videos at each of the above two resolutions to fit better previous methods. For each method, we resize the video to its default resolution before editing, then resize the edited video back to the original video resolution.
@@ 5.2 Quantitative Results
We first make a comprehensive quantitative comparison between our InsViE model and state-of-the-art video editing methods. As shown in Table 4, we evaluate the edited videos in three aspects with a total of eight metrics. One can see that our method achieves the best results in all metrics. In terms of temporal consistency, our method achieves a CLIP score of 0.956, indicating its superior ability to maintain coherence across video frames. Meanwhile, InsViE attains the lowest optical flow EPE, confirming its effectiveness in preserving motion dynamics and seamless transitions. Notably, the trend of GPT temporal score is not entirely consistent with EPE, indicating that GPT-4o cannot fully capture the video motion dynamics using the partially sampled frames. Therefore, incorporating optical flow into the filtering process is essential. In terms of textual alignment, InsViE achieves significantly higher CLIP score than its competitors, demonstrating strong capability in aligning visual content with textual descriptions. The Pick score and GPT score show similar trend, further underscoring the accuracy of InsViE in generating contextually appropriate content. Finally, InsViE surpasses previous methods on GPT quality score and DOVER score, confirming the high visual quality of produced videos. Overall, these quantitative results demonstrate InsViE’s comprehensive advantages in instruction-based editing over existing methods.
@@ 5.4 Ablation Study
Table 5: Ablation study on data filtering, multi-stage training and static-real ratio in the final training stage, where OF denotes optical flow.
@@
Static:Real=0:1 0.951 4.88 3.82 19.15 18.70 3.79 0.519 3.65
Static:Real=0.5:1 0.954 4.89 3.86 19.21 18.73 3.80 0.540 3.71
Static:Real=1:1 0.956 4.85 3.87 19.18 18.69 3.79 0.547 3.72
Static:Real=5:1 0.956 4.84 3.87 19.37 18.91 3.84 0.567 3.79
Ablation for data filtering. In Tab. 5, we investigate the role of GPT-4o [1] and Optical Flow [35] (OF) filters. Removing the GPT-4o filter will decrease the performance in all three aspects, especially the textual alignment, indicating a significant loss of coherence with textual descriptions. Similarly, excluding the OF filter decreases all metrics while producing the most significant negative impact on temporal consistency. Therefore, both filters play crucial roles in maintaining coherence and motion dynamics.
Ablation for multi-stage training. We then evaluate the effects of multi-stage training strategy in Tab. 5 and Fig. 6. In “Stage 1”, using Set-S1 yields competitive performance, while the following training stages provide further improvements. Specifically, “Stage 1&2” brings considerable gain of OF EPE and textual alignment. In  Fig. 6, “Stage 1&2” is better aligned with the source video with more stylistic elements. “Stage 1&2&3” brings notable improvements in textual alignment and video quality. Compared with “Stage 1&2”, the DOVER and Pick scores increase favorably by 0.052 and 0.26, illustrating the benefits of LPIPS loss and high-quality static videos. The example videos in Fig. 6 support the quantitative comparison. In short, the results underscore the necessity of each training stage.
@@
Ablation for static-real ratio. We further investigate the impact of different ratios of static to real videos in Set-S3. In Tab. 5, “Static:Real=0:1” exhibits similar results to “Stage 1&2”, indicating the limitation of using real videos only. As shown in Fig. 6, the result is smooth and lacks details. Increasing the ratio to “0.5:1” leads to better results than “Stage 1&2” on all the metrics. By setting “Static:Real=1:1”, the model’s performance stabilizes with better DOVER and GPT quality scores, demonstrating the benefits of static videos for visual quality. The most notable gain can be observed at “Static:Real=5:1”, especially on the textual alignment and video quality. In Fig. 6, “S:R=5:1” provides the best alignment with the source video and instruction, offering the most satisfactory visual quality. This ratio effectively incorporates the advantages of static and real videos, leading to the best editing ability. Thus, the 5:1 ratio was determined to be our optimal choice.
@@ 6 Conclusion
Limitations. Our InsViE-1M dataset and trained InsViE model have some limitations. First, our elaborately designed filtering process, though effective, is constrained by the visual understanding capability of GPT-4o. Second, the video clips in our dataset, while longer than existing video editing datasets, are still not long enough, which may limit the performance of trained models in processing longer videos. Lastly, our InsViE model is fine-tuned on open-source video generation models, and hence may inherit their limitations in synthesizing realistic videos.
@@
In this supplementary file, we provide additional details of the construction pipeline of our InsViE-1M dataset in Sec. 7, additional settings of model training and testing in Sec. 8, and more visual comparisons in Sec. 9. In addition, we provide a demo video that includes more visual comparisons. Please view the video using software that can open MOV files.

### v2

@@ 1 Introduction
However, the training data construction for instruction-based video editing (InsViE) is much more challenging than that for image editing. InstructVid2Vid [27] and EffiVED [42] create edited videos utilizing one-shot methods [35, 23], which is highly time consuming. As a result, only 24K videos with 8 frames per video of 256$\times$256 resolution are provided in  [42], whose video quantity and quality are not sufficient to train a robust video editor. InsV2V [6] synthesizes 400K editing pairs; however, the source videos are synthetic, and hence the trained models are less effective when editing real-world videos. In addition, the limited video resolution (256$\times$256) of this dataset limits the application of the trained models. As summarized in Tab. 1, existing InsViE datasets suffer from low-resolution, short video duration, and insufficient quantity and quality, making InsViE remain a challenging task.
@@ 2 Related Work
Training-free video editing. Many existing video editing methods rely on pre-trained large image/video models [26, 10, 20, 17, 8, 7, 18, 9] to perform DDIM [31] inversion and denoising without training on any data. FateZero [26] introduces attention blending to disentangle the editing of objects and background. To ensure motion consistency, FLATTEN [8] and RAVE [17] use optical flow and random noise shuffling to improve the editing effect. Tokenflow [10] and VidToMe [20] merge the intermediate tokens to accelerate the editing process. AnyV2V [18] and Videoshop [9] adopt pre-trained image-to-video (I2V) models [4, 41] to edit the whole video with a given edited first-frame, providing more interaction with users. However, training-free methods often yield unsatisfactory results due to the inherent gap between generation and editing tasks. Furthermore, it is difficult to ensure the quality of the generated samples when we construct large-scale datasets directly using training-free methods. Therefore, we design an elaborate pipeline to filter both the edited first frame and the final edited video, which significantly improves the quality of our dataset.
@@
Training-based video editing. Recent works [5, 6, 42, 15, 21] have started to build large video editing datasets to train video editing models. ViViD-10M [15] and GenProp [21] utilize mask and the first frame to edit the video by pre-trained large models. As initiated by the work of InstructVid2Vid [27], a class of methods [5, 6, 42] adopt the instruction-based editing approach. For more robust performance, InsV2V [6] creates a large dataset with more than 400k synthetic pairs and trains the editing model by introducing temporal layers to the model of InstructPix2Pix [5]. Since image editing models often fail to preserve temporal consistency, EffiVED [42] generates 155k training pairs, including 24k real-world source videos, to train a video editing model upon video generation model. Senorita-2M [44] builds multiple expert models to create high quality editing effects, creating 2M samples. However, fine-tuning a pre-trained video generation model requires extensive high-quality triplets (source video, edited video, instruction) to achieve sufficient instruction-based editing capability. In this paper, we construct a large video editing dataset for training instruction-based editing models by collecting high-quality source videos and designing effective filtering strategies to improve training data quality. Compared with Senorita-2M that uses trained classifier and CLIP for filtering, we apply multi-stage filtering by GPT-4o. Actually, these strategies can be integrated to further improve the quality of data.
@@ 3.1 Triplet generation from real-world videos
We curate high-quality videos from Pexel [2] and Openvid [22], and design an elaborate two-stage pipeline to generate and screen the edited samples. As shown in Fig. 2, we perform first frame editing and screening in the $1^{st}$ stage, and video editing and filtering in the $2^{nd}$ stage.
@@ 3.2 Triplet generation from image editing pairs
Existing image editing datasets [5, 40] possess a variety of image pairs with diverse editing types, which can be employed to generate video editing data. We first select 150k image editing triplets from InstructPix2Pix [5] based on metrics such as CLIP similarity, CLIP directional score, and the GPT-4o score we used in Sec. 3.1. We then employ all the 10k image editing pairs in MagicBrush [40] since this dataset is manually annotated with precise editing effects. As shown in Fig. 3, with these image editing pairs, we first generate the source videos from the source images through I2V generation model SVD [4]. Then, we perform video editing and filtering using the same process as the $2^{nd}$ stage used in Sec. 3.1. Finally, we generate 110k filtered video editing triplets from image editing pairs.
@@ 3.3 Generate static video triplets from images
Since producing video editing data is costly, previous works [15, 42, 33] have proposed to utilize images as part of data for mixed training. However, directly training on image data may not align well with real-world videos, hindering the training of video editing models. Considering the fact that a portion of real-world videos feature only on camera movement, we translate the images into static videos with manual camera operations, including (1) zoom-in, (2) zoom-out, (3) move left, (4) move right, (5) move down, (6) move up. The edited image is generated by using CosXL [3] as in the $1^{st}$ stage of Sec. 3.1. We crop each source and edited image pair at equal intervals to create image sequence pair, which is then processed with bilinear interpolation to enhance the video smoothness. Then we apply the filtering process as in the $2^{nd}$ stage of Sec. 3.1. In total, we produce 450k triplets of static videos that only possess camera movement for training.
@@ 4 InsViE Model Training
$\mathcal{L}$ $=\mathbb{E}_{\epsilon\sim\mathcal{N}(0,1),t}\left[\|\epsilon-\epsilon_{\theta}(t,z_{t},z_{source},z_{text})\|_{2}^{2}\right],\vspace{-0.2cm}$ (1)
@@ 4.2 Training and sampling
$\quad+s_{T}\cdot\left(\epsilon_{\theta}(z_{t},z_{source},z_{text})-\epsilon_{\theta}(z_{t},\varnothing,\varnothing)\right)\vspace{-0.15cm}$
@@ 5.1 Experimental Settings
Implement details. Our InsViE model is finetuned on the CogVideoX-2B model [39] using LoRA [14] with rank of 128. The training resolution is set to $720\times 480$ by random cropping on the $1024\times 576$ resolution source/edited videos in our InsViE-1M dataset. The training is conducted on 64 NVIDIA A100 GPUs for 40k iterations. The batch-size is 128. To be specific, we train our model on Set-S1, Set-S2 and Set-S3 for 20k, 10k, 10k iterations, respectively. Overall, it takes about 90s/20G/1GPU and 100h/75G/64GPUs for inference and training.
Compared methods. We compare our model with several representative methods, including SD-based training-free methods FateZero [26], TokenFlow [10] and RAVE [17], SVD-based training-free method Videoshop [9], and training-based method InsV2V [6], and EVE [30].
@@
[…]
Testing dataset. As in previous methods [10, 9, 17], we collect 100 video samples from DAVIS [25], YoutubeVOS [38] and Pexels [2] as the test set. We caption the videos to generate the corresponding instructions and paired captions using large VLMs [13] for instruction-based and training-free methods, respectively. Considering the fact that our method can process $720\times 480$ resolution video, while the competing methods can only process $512\times 512$ or lower resolution video, we collect 50 videos at each of the above two resolutions to fit better previous methods. For each method, we resize the video to its default resolution before editing, then resize the edited video back to the original video resolution. We also compare our model with baselines on TGVE [36] and TGVE+ [30]. We conduct experiments by resizing the videos to $720\times 480$ and resizing back to $480\times 480$ after editing for evaluation.
@@ 5.2 Quantitative Results
We first make a comprehensive quantitative comparison between our InsViE model and state-of-the-art video editing methods. As shown in Tab. 4, we evaluate the edited videos in three aspects with a total of eight metrics. One can see that our method achieves the best results in all metrics. In terms of temporal consistency, our method achieves a CLIP score of 0.956, indicating its superior ability to maintain coherence across video frames. Meanwhile, InsViE attains the lowest optical flow EPE, confirming its effectiveness in preserving motion dynamics and seamless transitions. Notably, the trend of GPT temporal score is not entirely consistent with EPE, indicating that GPT-4o cannot fully capture the video motion dynamics using the partially sampled frames. Therefore, incorporating optical flow into the filtering process is essential. In terms of textual alignment, InsViE achieves significantly higher CLIP score than its competitors, demonstrating strong capability in aligning visual content with textual descriptions. The Pick score and GPT score show similar trend, further underscoring the accuracy of InsViE in generating contextually appropriate content. Finally, InsViE surpasses previous methods on GPT quality score and DOVER score, confirming the high visual quality of produced videos. We can see similar trends on the TGVE/TGVE+ datasets in Tab. 6 . We compare our method with TokenFlow, InsV2V and EVE, which are the best performers on TGVE. Overall, these quantitative results demonstrate InsViE’s comprehensive advantages in instruction-based editing over existing methods.
@@ 5.4 Ablation Study
Table 5: Ablation study on data filtering and multi-stage training, where OF denotes optical flow.
@@
Table 6: Quantitative comparison on TGVE/TGVE+. The best results are highlighted in bold. indicates retraining on our dataset.
[…]
TokenFlow [10] 20.58/20.62 0.943/0.944 0.117/0.085 0.257/0.254
InsV2V [6] 20.76/20.37 0.911/0.925 0.208/0.174 0.262/0.236
EVE [30] 20.76/20.88 0.922/0.926 0.221/0.198 0.262/0.251
InsV2V [6] 20.83/20.91 0.931/0.939 0.237/0.191 0.271/0.252
Ours 20.90/20.97 0.941/0.945 0.242/0.201 0.278/0.270
Ablation on data filtering. In Tab. 8, we investigate the role of GPT-4o [1] and Optical Flow [37] (OF) filters. Removing the GPT-4o filter will decrease the performance in all three aspects, especially the textual alignment, indicating a significant loss of coherence with textual descriptions. Similarly, excluding the OF filter decreases all metrics while producing the most significant negative impact on temporal consistency. Therefore, both filters play crucial roles in maintaining coherence and motion dynamics.
Ablation on multi-stage training. We then evaluate the effects of multi-stage training strategy in Tab. 8 and Fig. 6. In “Stage 1”, using Set-S1 yields competitive performance, while the following training stages provide further improvements. Specifically, “Stage 1&2” brings considerable gain of OF EPE and textual alignment. In  Fig. 6, “Stage 1&2” is better aligned with the source video with more stylistic elements. “Stage 1&2&3” brings notable improvements in textual alignment and video quality. Compared with “Stage 1&2”, the DOVER and Pick scores increase favorably by 0.052 and 0.26, illustrating the benefits of LPIPS loss and high-quality static videos. The example videos in Fig. 6 support the quantitative comparison. In short, the results underscore the necessity of each training stage.
Ablation on static-real ratio. In Fig. 6, “S:R” denotes the ratio between static and real videos. “S:R=5:1” provides the best alignment with the source video and instruction, offering the most satisfactory visual quality. This ratio effectively incorporates the advantages of static and real videos, leading to the best editing ability. Thus, the 5:1 ratio was determined to be our optimal choice.
@@
Ablation on the pre-trained baseline model. In Tab. 6, we train InsV2V on our dataset (InsV2V). The improvements demonstrate the contribution of our dataset. More ablation studies can be found in the supplementary file.
@@ 6 Conclusion
Limitations. Our InsViE-1M dataset and trained InsViE model have some limitations. First, our elaborately designed filtering process, though effective, is constrained by the visual understanding capability of GPT-4o. Second, the video clips in our dataset, while longer than existing video editing datasets, are still not long enough, which may limit the performance of trained models in processing longer videos. Third, our model performs slightly worse when adding useless content or descriptions to the input instructions because our training set includes mainly clean data. Lastly, our InsViE model is fine-tuned on open-source video generation models, and hence may inherit their limitations in synthesizing realistic videos.
@@
In this supplementary file, we provide additional details of the construction pipeline of our InsViE-1M dataset in Sec. 7, additional settings of model training and testing in Sec. 8, more visual comparisons in Sec. 9, and more ablation studies in Sec. 10. In addition, we provide a demo video that includes more visual comparisons. Please view the video using software that can open MOV files.
@@ 9 More Visual Results
Table 8: Ablation study on static-real ratio in the final training stage.
Static:Real=0:1 0.951 4.88 3.82 19.15 18.70 3.79 0.519 3.65
Static:Real=0.5:1 0.954 4.89 3.86 19.21 18.73 3.80 0.540 3.71
Static:Real=1:1 0.956 4.85 3.87 19.18 18.69 3.79 0.547 3.72
Static:Real=5:1 0.956 4.84 3.87 19.37 18.91 3.84 0.567 3.79
10 More Ablation Studies
Table 9: Ablation study on the LPIPS loss in Stage 3.
[…]
w/ LPIPS 3.88 3.84 0.567 3.79
w/o LPIPS 3.86 3.83 0.543 3.73
Ablation on the LPIPS loss in Stage 3. As described in Sec. 4.2 in the main paper, we use $L_{2}$ loss in the first two training stages. LPIPS loss is added in the final stage to enhance detail generation. As shown in Tab. 9, it contributes more to video quality metrics.
Ablation on static-real ratio. We further investigate the impact of different ratios of static to real videos in Set-S3. In Tab. 8, “Static:Real=0:1” exhibits similar results to “Stage 1&2”, indicating the limitation of using real videos only. Increasing the ratio to “0.5:1” leads to better results than “Stage 1&2” on all the metrics. By setting “Static:Real=1:1”, the model’s performance stabilizes with better DOVER and GPT quality scores, demonstrating the benefits of static videos for visual quality. The most notable gain can be observed at “Static:Real=5:1”, especially on the textual alignment and video quality.